OpenStudy (anonymous):

Solve negative 5 over 2, the whole multiplied by x minus 5 equals negative 55.

Mehek (mehek14):

\(\dfrac{5}{2}x-5=-55\) ?

OpenStudy (anonymous):

a. -20 b. -24 c. 20 d. 24

Mehek (mehek14):

does the equation look like what I put?

OpenStudy (anonymous):

(-5/2)x - 5 = -55

OpenStudy (anonymous):

what is x

Mehek (mehek14):

\(-\dfrac{5}{2}x-5=-55\) ?

OpenStudy (anonymous):

no only the 5 is negitve.

OpenStudy (serenity74):

Solve for X. Bring 5 to the other side.

Mehek (mehek14):

it's the same thing \(-\dfrac{5}{2}x-5=-55\) \(\dfrac{-5}{2}x-5=-55\)

OpenStudy (anonymous):

the second on.

Mehek (mehek14):

first thing -55 + 5 = ?

OpenStudy (serenity74):

Then in order to remove the fraction you multiply by the reciprocal.

OpenStudy (serenity74):

and bam you get the answer for X.

OpenStudy (danjs):

the 'whole' , quantity possibly

OpenStudy (anonymous):

what is x?

OpenStudy (serenity74):

You would get a whole number even if you did the reciprocal....

Mehek (mehek14):

first -55 + 5 = ?

OpenStudy (anonymous):

-50

OpenStudy (danjs):

(-5/2) * ( x - 2 ) = -55 maybe

Mehek (mehek14):

\(\dfrac{-5}{2}=-50\)

OpenStudy (serenity74):

Wat are you doing @DanJS lol

Mehek (mehek14):

\(\dfrac{2}{-5}*-50\)

OpenStudy (serenity74):

The X is part of 5/2 it's not seperated lol.

OpenStudy (anonymous):

20

OpenStudy (serenity74):

Ummmmm @Mehek14 dats a bit too complicated

Mehek (mehek14):

that's how you solve it @Serenity74

OpenStudy (serenity74):

I lied my bad

OpenStudy (serenity74):

I didn't that you reciprocal-ed it sorry!!!!

OpenStudy (anonymous):

so the answer is 20?

Mehek (mehek14):

yup

OpenStudy (serenity74):

Yes the answer is 20 and @Mehek14 is correct.

OpenStudy (anonymous):

thank you!

Mehek (mehek14):

2*-50 = -100 -100/-5 = +20

OpenStudy (danjs):

(-5/2) * ( x - 5 ) = -55 sorry the wording looked like that to me, multiplied by (x - 5), that gives x=27 though

OpenStudy (anonymous):

thank you!!

OpenStudy (serenity74):

Wait.

OpenStudy (anonymous):

yes

OpenStudy (serenity74):

No it is 20

OpenStudy (danjs):

instead of multiplied by x - and take away a 5 , or something

OpenStudy (anonymous):

@DanJS 27 is not an answer.

OpenStudy (danjs):

yeah , sorry

OpenStudy (anonymous):

@Serenity74 thank you.

OpenStudy (anonymous):

good bye and I might have more questions.

OpenStudy (serenity74):

Yes so you end up with -5/2x = -50 -5x = -50 (2) (since the two cancel out on the other side by multiplying it) then you get -5x = -100 x = -100/-5 = +20 (since negative * negative give you a positive answer) @DanJS

OpenStudy (serenity74):

Yes so you end up with -5/2x = -50 -5x = -50 (2) (since the two cancel out on the other side by multiplying it) then you get -5x = -100 x = -100/-5 = +20 (since negative * negative give you a positive answer) @DanJS

OpenStudy (danjs):

right, the wording in the problem is goofy to me if the words had something like 'and take aqway 5' or 'then take 5 away', may be clearer the x - 5 sounds like a quantity to me, sorry bout that

OpenStudy (danjs):

multiplied by x minus 5 -- to me is (x - 5)

OpenStudy (serenity74):

Lol but you do get it right now right?

OpenStudy (danjs):

yes i know how to do the math, just the wording is different than i have seen, maybe

OpenStudy (serenity74):

Lol you don't have to be sorry about nothing lol, wordings tend to trip people off XD.

OpenStudy (serenity74):

Heyyyyyyy IK you know math lol you after all have 99 and I am just at 41 lol. No needa get defensive lol.

OpenStudy (serenity74):

Heyyyyyyy IK you know math lol you after all have 99 and I am just at 41 lol. No needa get defensive lol.

OpenStudy (serenity74):

Heyyyyyyy IK you know math lol you after all have 99 and I am just at 41 lol. No needa get defensive lol.

OpenStudy (danjs):

are you sure there isnt something between the 'X' and the 'minus' like even a comma, or something

OpenStudy (serenity74):

Heyyyyyyy IK you know math lol you after all have 99 and I am just at 41 lol. No needa get defensive lol.

OpenStudy (serenity74):

Heyyyyyyy IK you know math lol you after all have 99 and I am just at 41 lol. No needa get defensive lol.

OpenStudy (danjs):

haha , you can get to 99, with medals from addition probs from 1st grade maybe in a long long time

OpenStudy (serenity74):

Heyyyyyyy IK you know math lol you after all have 99 and I am just at 41 lol. No needa get defensive lol.

OpenStudy (serenity74):

XD @DanJS I am here to get some Calculus practice actually but i DON'T SEE ANY CALCULUS!! GAH! XD

OpenStudy (serenity74):

XD @DanJS I am here to get some Calculus practice actually but i DON'T SEE ANY CALCULUS!! GAH! XD

OpenStudy (danjs):

umm

OpenStudy (danjs):

some of those are a lot of pages of prob solutions

OpenStudy (danjs):

to learn some tools to look for in your own probs

OpenStudy (serenity74):

Thanks! @DanJS

OpenStudy (serenity74):

Thanks! @DanJS